Qn 4. For a given angle θ, suppose that

sin2 θ + cos2 θ = 1 and cos2 θ − sin2 θ = cos 2θ.

2Derive an expression for cos 4θ in terms only of powers of cos θ


1
Expert's answer
2022-02-28T16:36:40-0500
"\cos 2θ=\cos ^2θ-\sin ^2θ=2\cos^2\theta-1"

"\cos 4θ=2\cos^22\theta-1=2(2\cos^2\theta-1)^2-1"

"=8\cos^4\theta-8\cos^2\theta+1"